Do the data represent a direct variation or an inverse variation? Write an equation to model the date.
a. (-2, 0.16), (-10, 0.8), (50, -4)
b. (0.5,100), (2.5,20),(10,5)


Please do both!!!! Its a two parter!
An equation for both please :)

Answer by Boreal(15235)

b is easier
It is inverse variation because as y increases, x decreases or vice versa.
y=k/x. Find k. Use 100=k/0.5; k=50; check in other two 20=k/2.5=50/2.5, check; 5=k/10=50/10 checks.
y=50/x
------------------------------
a. As x goes from -10 to 50, y goes from 0.8 to -4, there is a direct variation with what looks like a factor of -5
y=kx
-4=50k
k=-4/50=-2/25
y=-0.08/x
=================
check in other two.
0.8=(-2/25)(-10)=20/25, checks
0.16=(-2/25)(-2), checks.